OpenStudy (anonymous):

PLEASE I NEED HELP ASAP Andrew is showing his work in simplifying −4.5 + 4.2 + 5.6 − 7.3. Identify any errors in his work or in his reasoning. Write feedback to Andrew explaining the mistake he made and how to correct his work. Step Work Reasoning −4.5 + 4.2 + 5.6 − 7.3 Original problem 1 −4.5 + 4.2 + 5.6 + (−7.3) Additive inverse 2 −4.5 + (−7.3) + 4.2 + 5.6 Commutative property 3 (−4.5 + (−7.3)) + (4.2 + 5.6) Associative property 4 −11.8 + 9.8 Simplify 5 −21.6 Simplify

OpenStudy (anonymous):

−4.5+4.2+5.6−7.3 =−0.3+5.6−7.3 =5.3−7.3 =−2

OpenStudy (anonymous):

So whats his error

OpenStudy (anonymous):